Pls help me solve this problem...."the perimeter of a rectangle is 10 meters , and twice the width is equal to 1/2 the length. Find the dimension of the rectangle? "
Thank u ^^

Answer by macston(5194)

.
This is the last time I will respond to a post with abbreviations such as "pls" and "u". If you cannot go to the trouble of typing out your post, I do not have the time to respond.
.
P=perimeter; L=length; W=width
.
2W=1/2L
4W=L
.
P=2(L+W)
10m=2(4W+W)
10m=10W
1m=W ANSWER 1 The width is 1 meter.
L=4W
L=4(1m)=4m ANSWER 2: the length is 4 meters.